a non-uniform rod xy of mass 40kg and length 5m lies on the horizontal ground.its center of gravity is 2m from the end x. calculate the vertical force p, that will just be sufficient to lift the end y from the ground. why would the force p, if applied at the end x not be sufficient to lift the end of x from the ground?( g=10ms-2).


1
Expert's answer
2021-05-12T17:55:33-0400

Taking the moments about the end X:


"PL−mg(d)=0\\
P(5)-(40)(9.8)(2)=0\\
P=157\ N"

If force P was applied at the end X. Taking the moments about the end Y:


"M=PL−mg(L−d)\\
M=157(5)-(40)(9.8)(5-2)=-391<0"